OpenStudy (anonymous):

i want to ask the cube of 3x-1/3y plzz answer this algebraic question with steps

OpenStudy (anonymous):

There is an algebraic identity for such question. You want to cube of (3x-1/3y). Such type is identified by (a-b)^3. i.e \[a ^{3}-b ^{3}-3ab(a-b)\] Now put the value of a=3x and b=1/3y and solve it.

OpenStudy (anonymous):

ACTUALLY THATS WHAT I CANNOT DO ALTHOUGH I KNOW THE IDENTITY PLZZ HELP ME WITH THE STEPS

OpenStudy (anonymous):

where do you get stuck

OpenStudy (anonymous):

FROM THE FIRST STEP

OpenStudy (anonymous):

see since a=3x and b=1/3y, first find the cube of both a and b i.e a^3=27x^3 and b^3=1/27y^3. Do you get it...

OpenStudy (anonymous):

WHAT DOES ^ THIS SIGN MEANS

OpenStudy (anonymous):

raise to power of

OpenStudy (anonymous):

ok yah am getting it then

OpenStudy (anonymous):

now calculate the value of 3a^2b i.e \[3\times(3x)^{2}\times(1/3y)\] Which is equal to \[9x ^{2}/y\]

OpenStudy (anonymous):

are you there